# Mitigate the Gap: Investigating Approaches for Improving Cross-Modal Alignment in CLIP

### v2

@@ 4.3 Zero-Shot Transfer
CLIP’s pre-training objective for predicting whether a text is paired with an image has resulted in outstanding zero-shot image classification and multi-modal retrieval results (Radford et al., 2021). Therefore, we further assess how sharing the learnable parameters and the intra-modality loss affect the zero-shot transfer to donstream tasks. Furthermore, Section A.2 reports the results for image classification with linear probing.
@@
Image Classification. We conduct the zero-shot classification experiments on ImageNet-1K (Russakovsky et al., 2015), CIFAR-100, CIFAR-10 (Krizhevsky et al., 2009), Food-101 (Bossard et al., 2014), and Stanford Cars (Krause et al., 2013) with the combination of text prompts used by CLIP (Radford et al., 2021), e.g., “a photo of the {label}”. The experimental results summarized in Table 2 show that SharedCLIP reduces the accuracy of the zero-shot classification on CIFAR-10 by about $5\%$ and $0.4\%$ when measuring Top-1 and Top-5 accuracy scores, respectively. Similarly, SharedCLIP’s results on CIFAR-100 show about $2\%$ and $1\%$ decrease in accuracy. The trend of accuracy reduction is also observed on Food-101 and Stanford Cars datasets. However, on ImageNet-1K, SharedCLIP evinces up to $1\%$ improvement of accuracy. In contrast, AlignCLIP yields the best scores across all five datasets. On CIFAR-10, AlignCLIP achieves up to $5\%$ and $1\%$ improvements in Top-1 and Top-5 accuracy scores, when compared to the original CLIP model. Similarly, using AlignCLIP for CIFAR-100 results in about $2\%$ and $6\%$ enhancement in Top-1 and Top-5 accuracy scores in comparison to the original CLIP. The trend of improvement is also observed on the Stanford Cars dataset. Furthermore, AlignCLIP achieves $1\%$ and $1.3\%$ improvement of accuracy on ImageNet-1K when compared to the original CLIP. On the Food-101 dataset, the original CLIP model achieves the best Top-1 accuracy of 35.2 and AlignCLIP achieves comparable results in Top-5 accuracy. According to our ablation study, AlignCLIP without sharing the parameters between encoders results in a noticeable decrease of the classification performance. Our experiments thus evince that via sharing parameters and the additional intra-modality separation, AlignCLIP improves the cross-modal alignment of the embeddings without losing performance on the downstream zero-shot image classification task.
@@
Multi-Modal Retrieval. In addition to classification, we evaluate SharedCLIP and AlignCLIP in the application of zero-shot image-to-text and text-to-image retrieval using the validation splits from the MSCOCO (Lin et al., 2014) and Flickr30K (Plummer et al., 2015) datasets. The results of these evaluations are reported in Table 3. In all settings, the text prompt “a photo of the {caption}” is used. Our experiments show that both SharedCLIP and AlignCLIP improve the retrieval results measured by $R@\{1,5,10\}$ on both datasets when compared to the original CLIP model. In addition, AlignCLIP achieves comparable results to SharedCLIP in $R@\{5,10\}$. In summary, the best results for image-to-text and text-to-image retrieval on both MSCOCO and Flickr30K datasets are achieved by SharedCLIP, with AlignCLIP ranking in a very close second position. Finally, according to our ablation, AlignCLIP without sharing the parameters between encoders decreases the retrieval performance in comparison to the original CLIP model. These experiments demonstrate that it is possible to reduce the modality gap in CLIP via parameter sharing while improving the downstream multi-modal retrieval tasks. Furthermore, the addition intra-modality separation improves the alignment noticeably without losing perofamnce when compared to the original CLIP model.
@@
Semantic Text Similarity. We further evaluate SharedCLIP and AlignCLIP models on seven Semantic Textual Similarity (STS) benchmarks using the SemEval toolkit (Conneau and Kiela, 2018). Measuring the level of semantic similarity between encoded texts using cosine similarity, we report the Pearson correlations between the human-assigned similarity scores, provided as the ground-truth in the STS benchmark datasets, and the cosine similarity of the encoded texts in Table 4. SharedCLIP shows slight improvements on STS14 and STS15 while reducing the correlation on the rest of the benchmarks. This shows that sharing the transformer between the image and text modalities damages the semantic representation of the texts. In contrast, the intra-modality separation with the semantic similarity re-scaling mechanism in AlignCLIP substantially corrects this damage and improves the results noticeably. As a result, AlignCLIP achieves the best average results on the STS task. Similar to the previous tasks, AlignCLIP without sharing parameters substantially decreases the performance. These observations consistently show that AlignCLIP, while reducing the modality gap, achieves comparable results to CLIP in the zero-shot evaluation of semantic text similarity.
Model
STS12
STS13
STS14
STS15
STS16
STS-B
SICK-R
Average
CLIP 0.55 0.42 0.54 0.59 0.60 0.59 0.79 0.58
SharedCLIP 0.55 0.40 0.55 0.60 0.58 0.59 0.72 0.57
AlignCLIP (no sharing) 0.52 0.36 0.53 0.58 0.60 0.58 0.78 0.56
AlignCLIP 0.56 0.43 0.60 0.64 0.64 0.59 0.77 0.6
Table 4: Pearson correlations for zero-shot STS using the text embeddings.
4.4 Robustness to Natural Distribution Shift
In the zero-shot image classification task, CLIP has additionally shown impressive robustness to natural distribution shifts and promising generalizability to out-of-distribution images. Therefore, we expand our evaluations and investigate to what extent SharedCLIP and AlignCLIP change the performance with natural distribution shifts. We use the ImageNetV2, ImageNet-R, ImageNet-A, ImageNet-O and ImageNetSketch datasets for these evaluations and report the corresponding results in Table 5, in terms of Top-1 and Top-5 accuracy. In summary, SharedCLIP achieves the best results on ImageNet-R, ImageNet-O and ImageNetSketch, whereas the best results on ImageNetV2 and ImageNet-A are achieved by AlignCLIP. Based on these observations, we conclude that sharing parameters and applying intra-modality separation in SharedCLIP and AlignCLIP preserve the robustness to natural distribution shifts when compared to the original CLIP model and at the same time, improve the cross-modal alignment. However, the IMSep objective in AlignCLIP without sharing parameters of the encoders results in a decrease in the classification accuracy.
Model ImageNetV2 ImageNet-R ImageNet-A ImageNet-O ImageNetSketch
Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
CLIP 27.1 53.3 39.8 65.8 6.5 25.4 25.8 45.5 19.4 41.8
SharedCLIP 27.5 53.5 40.2 67.3 6.7 25.5 26.2 45.8 20.6 43.2
AlignCLIP (no sharing) 24.6 50 34.5 59.7 5.8 23.2 23.2 42 17.6 38.6
AlignCLIP 27.7 53.9 38.6 65 6.9 25.5 24.8 45.8 19.7 42
4.5 DOSNES Visualizations
@@
For a more comprehensive comparison of the distribution of each modality, we visualize the DOSNES projection of the encoded image–texts from the MSCOCO validation set in Figure 5. As can be seen, the uni-modal embeddings in CLIP are densely located on opposite sides of the hypersphere. In contrast, the embeddings get spread out when using the SharedCLIP model. Finally, the embedding space of AlignCLIP achieves the best spread of the uni-modal embeddings and substantially greater alignment of image and text embeddings. Thus, the intra-modality separation leads to a better alignment and a substantial reduction of the modality gap. Furthermore, Figure 5 shows that AlignCLIP reduces the sparsity of the embeddings on the hypersphere.
Figure 5: DOSNES visualization of the multi-modal embeddings using CC3M.
@@ 5 Related Work
Modality Gap in CLIP. The modality gap in CLIP’s embedding space was first studied by Liang et al. (2022). Authors showed that the modality gap is caused by a combination of the model initialization and the contrastive loss optimization. Furthermore, they showed that CLIP’s embedding space is very sparse such that the effective embedding space is an extremely narrow cone. In a similar study, Tyshchuk et al. (2023) measured the alignment of image and text embeddings in CLIP with a focus on the isotropic properties. More recently, Schrodi et al. (2024) showed that one of the key factors contributing in both the modality gap and the bias in CLIP is the information-imbalance between the two modalities, i.e., images often have much more detailed information of the scene in comparison to their corresponding text descriptions. Furthermore, Jiang et al. (2023) showed that exact modality alignment in multi-modal embeddings is suboptimal for downstream tasks and propose intra-modality as well as inter-modality regularization using augmented samples in order to improve the latent embeddings structures. In contrast to the previous work, we study the effectiveness of parameter sharing between the modality encoders as well as semantically-regularized separation of the uni-modal embeddings on reducing the modality gap.
@@ 6 Conclusion
This work investigates the potential of reducing the modality gap in CLIP by sharing the learnable parameter space of the vision and language encoders as well as enforcing a contrastive intra-modality separation objective. We further examine the effects of modality gap reduction by the aforementioned refinements in the performance of downstream tasks. Through extensive experiments, we show that SharedCLIP and AlignCLIP improve various zero-shot downstream applications when compared to CLIP, while substantially improving the cross-modal alignment, and therefore, reducing the gap. Furthermore, our ablation on the intra-modality loss shows that sharing parameters between the vision and language encoders is necessary for preventing decrease in the performance of the downstream tasks. Our work shows that it is possible to mitigate the modality gap in CLIP via parameter sharing and intra-modality separation without losing downstream performance.
@@ A.1 Training Dataset and Setup
We adopted a transformer encoder consisting of 12 layers and 12 heads in SharedCLIP and AlignCLIP. The same input pre-processing and augmentations employed in CLIP have been used for SharedCLIP as well as AlignCLIP, including random cropping of images to the size $224\times 224$. The image patch size for encoding visual data is set to $16\times 16$. When encoding texts, the maximum sequence length is set to 77 tokens and the vocabulary size for the embedding layer is set to 49,408. The output embedding dimensionality for both the vision and language modalities is set to 768. We chose our setup to be quite similar to that of the original CLIP with the ViT-L-16 backend, in order to ensure a fair comparison of CLIP and our proposed model. SBERT all-mpnet-base-v2 model has been utilized in AlignCLIP for re-scaling text–text and image–image cosine similarities is the .
@@
A.2 Linear Probing for Image Classification
Table 6 shows the results of image classification with linear probing.
Model ImageNet-1K CIFAR-100 CIFAR-10 Food-101 Stanford Cars
CLIP 50 62.6 85 67 42.2
SharedCLIP 51.2 63 85 68.3 40.5
AlignCLIP 50.3 63.4 85.8 67.9 42.2
@@ A.2 Linear Probing for Image Classification
A.3 Experimental Results With Lower Output Dimensions
In order to examine how the trend of our observations change based on the choice of output embedding dimension, we repeated our training and downstream tasks using the output dimension of 512. We used the same hyperparatmer setting described in Section A.1. The corresponding results are reported in Table 7 to Table 9
Model CC3M MSCOCO ImageNet-1K CIFAR-100 CIFAR-10
CLIP 0.42 0.47 0.41 0.38 0.4
SharedCLIP 0.6 0.64 0.59 0.55 0.57
AlignCLIP 0.63 0.68 0.63 0.58 0.59
Table 7: Comparison of alignment scores when output dimension 512.
Model ImageNet-1K CIFAR-100 CIFAR-10 Food-101 Stanford Cars
Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
CLIP 31.5 59 27.6 57.1 65.6 95.6 33.2 61.9 11.6 37.4
SharedCLIP 31.6 59 29 58.7 59.5 95.6 33.7 62.5 9.4 32.7
AlignCLIP 30.9 58.6 26.5 53.7 53.3 92.8 31.7 60.5 9.7 33.4
Table 8: Accuracy scores for zero-shot image classification when output dimension 512.
MSCOCO Flickr30K
Model I $\rightarrow$ T T $\rightarrow$ I I $\rightarrow$ T T $\rightarrow$ I
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CLIP 31.7 57.9 68.9 20.5 44.3 56 54.5 78.7 87.3 40.3 68.6 78.8
SharedCLIP 32.8 59.3 70.4 21 44.6 56.6 56.1 81.5 87.3 41.2 69.5 78.9
AlignCLIP 30.6 56.8 68.5 20.3 43.6 55.4 54.6 80.5 88 40.6 68.1 77.5
Table 9: Zero-shot cross-modal retrieval summarized with R@{1, 5, 10} when output dimension 512.

### v3

@@ 4.3 Classification
CLIP’s pre-training objective for predicting whether a text is paired with an image has resulted in outstanding image classification performance when tested in a zero-shot setting as well as after linear probing (Radford et al., 2021). Therefore, we further assess how sharing the learnable parameters and the intra-modality loss affect the classification performance in these settings.
@@
Zero-Shot Image Classification. We conduct the zero-shot classification experiments on ImageNet-1K (Russakovsky et al., 2015), CIFAR-100, CIFAR-10 (Krizhevsky et al., 2009), Flowers-102 (Nilsback and Zisserman, 2008), and Stanford Cars (Krause et al., 2013) with the combination of text prompts used by CLIP (Radford et al., 2021), e.g., “a photo of the {label}”. The experimental results summarized in Table 2 show that SharedCLIP reduces the accuracy of the zero-shot classification on CIFAR-10 by about $5\%$ and $0.4\%$ when measuring Top-1 and Top-5 accuracy scores, respectively. Similarly, SharedCLIP’s results on CIFAR-100 show about $2\%$ and $1\%$ decrease in accuracy. The trend of accuracy reduction is also observed on Flowers-102 and Stanford Cars datasets. However, on ImageNet-1K, SharedCLIP evinces up to $1\%$ improvement of accuracy. In contrast, AlignCLIP yields the best scores across all five datasets. It achieves $1.4\%$ and $2\%$ improvement of Top1 and Top5 accuracy, respectively, on ImageNet-1K when compared to the original CLIP. On CIFAR-10, AlignCLIP achieves up to $8\%$ and $2\%$ improvements in Top-1 and Top-5 accuracy scores. Similarly, using AlignCLIP for CIFAR-100 results in about $8\%$ and $11\%$ enhancement in Top-1 and Top-5 accuracy scores in comparison to the original CLIP. The trend of improvement is also observed on the Flowers-102 and Stanford Cars datasets. Our experiments thus evince that via sharing parameters and the additional intra-modality separation, AlignCLIP improves the cross-modal alignment of the embeddings while enhancing the performance on the downstream zero-shot image classification task.
Linear Probing. We further test the performance of SharedCLIP and AlignCLIP when performing linear probing for image classification and report the Top1 accuracy results in Table 3. For all datasets, we train the linear classifier layer with a batch size of 128, for 30 epochs, with AdamW optimizer, and a cosine scheduler with a starting learning rate of 5e-4. Table 3 shows the superiority of AlignCLIP in the task of image classification with linear probing across all 5 datasets.
4.4 Robustness to Natural Distribution Shift
In the zero-shot image classification task, CLIP has additionally shown impressive robustness to natural distribution shifts and promising generalizability to out-of-distribution images. Therefore, we expand our evaluations and investigate to what extent SharedCLIP and AlignCLIP change the performance with natural distribution shifts. We use the ImageNetV2, ImageNet-R, ImageNet-A, and ImageNetSketch datasets for these evaluations and report the corresponding results in Table 4, in terms of Top-1 and Top-5 accuracy. It is first observed that SharedCLIP generally improves the classification accuracy in comparison to the CLIP model. Secondly, AlignCLIP achieves the best classification results across all datasets. In summary, when comparing to the CLIP model, AlignCLIP achieves about $2\%$ and $1\%$ improvement in Top1 and Top5 accuracy on the ImageNetV2 dataset. On ImageNet-R, AlignCLIP improves both Top1 and Top5 scores by about $2\%$. The positive general trend of the accuracy enhancement is also observed on the ImageNet-A and ImageNetSketch datasets. Based on these observations, we conclude that sharing parameters and applying intra-modality separation in SharedCLIP and AlignCLIP improves the robustness to natural distribution shifts when compared to the original CLIP model and at the same time, improve the modality gap.
Model ImageNetV2 ImageNet-R ImageNet-A ImageNetSketch
Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
CLIP 27.1 53.3 39.8 65.8 6.5 25.4 19.4 41.8
SharedCLIP 27.5 53.5 40.2 67.3 6.7 25.5 20.6 43.2
AlignCLIP 29.1 54.4 41.2 67.3 7 25.6 20.7 43.2
4.5 Multi-Modal Retrieval
@@
Zero-Shot Transfer. In addition to classification, we evaluate SharedCLIP and AlignCLIP in the application of zero-shot image-to-text and text-to-image retrieval using the validation splits from the MSCOCO (Lin et al., 2014) and Flickr30K (Plummer et al., 2015) datasets. The results of these evaluations are reported in Table 5. In all settings, the text prompt “a photo of the {caption}” is used. Our experiments show that both SharedCLIP and AlignCLIP improve the retrieval results measured by $R@\{1,5,10\}$ on both datasets when compared to the original CLIP model. In addition, AlignCLIP achieves the best overall results in comparison to SharedCLIP. When testing text-to-image retrieval on the MSCOCO dataset, SharedCLIP outperforms AlignCLIP at R@10. Similarly, for the image-to-text retrieval on the Flickr dataset, SharedCLIP achieves the best results. These experiments demonstrate that it is possible to reduce the modality gap in CLIP via parameter sharing while improving the downstream multi-modal retrieval tasks. Furthermore, the addition intra-modality separation improves the alignment noticeably while noticeably enhancing the retrieval performance.
@@
MSCOCO Flickr30K
Model I $\rightarrow$ T T $\rightarrow$ I I $\rightarrow$ T T $\rightarrow$ I
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CLIP 39.6 67.5 78.3 26.7 53.4 65.7 64.8 87.8 93.9 47.4 75.9 84
SharedCLIP 40.7 69.2 79.6 27.9 55 66.7 66.5 89.1 94.1 48.9 76.4 84.3
AlignCLIP 41.7 69.3 80.1 27.9 54.8 66.2 66 89.1 94.5 49.4 76.7 84.4
Table 6: Fine-tuned cross-modal retrieval summarized with R@{1, 5, 10}.
Fine-tuning Multi-Modal Retrieval. Table 6 summarizes the result of multi-modal retrieval when each model is fine-tuned using the corresponding training set. We fine-tuned each model for 8 and 20 epochs on MSCOCO and Flickr, respectively. In both cases, the batch size was set to 128 and the AdamW optimizer with the learning rate of 5e-6 and a weight decay of 0.2 was used. Our results show that both SharedCLIP and AlignCLIP outperform the CLIP model in the fine-tuning scenario. Additionally, AlignCLIP generally achieves a slightly better performance in comparison to SharedCLIP.
4.6 Ablation Study
This section provides an ablation study on the effectiveness of the re-scaling mechanism proposed in Eq. 9. In order to see the impacts on the different types of downstream tasks, i.e., image classification, classification with distribution shift and multi-modal retrieval, we compare the performance of AlignCLIP with and without the re-scaling mechanism on ImageNet-1K, CIFAR-100, CIFAR-10, ImageNetV2, MSCOCO and Flickr30K datasets and summarize the results in Table 7. This study substantially concludes that the re-scaling mechanism is effective in controlling the separation of similar image samples in the batch and therefore, increasing the results in the downstream tasks.
Model ImageNet-1K CIFAR-100 CIFAR-10 ImageNetV2 MSCOCO FLICKR
Top1 Top1 Top1 Top1 I $\rightarrow$ T T $\rightarrow$ I I $\rightarrow$ T T $\rightarrow$ I
AlignCLIP-W/O Rescaling 32.8 34.7 64.2 27.9 34 59.7 56 42.5
AlignCLIP 32.8 36.5 69.3 29.1 35.1 60.8 57.2 42.7
Table 7: Ablation study on the re-scaling mechanism in AlignCLIP.
4.7 Results Analysis
@@
DOSNES Visualization. For a more comprehensive comparison of the distribution of each modality, we visualize the DOSNES projection of the encoded image–texts from the MSCOCO validation set in Figure 5. As can be seen, the uni-modal embeddings in CLIP are densely located on opposite sides of the hypersphere. In contrast, the embeddings get spread out when using the SharedCLIP model. Finally, the embedding space of AlignCLIP achieves the best spread of the uni-modal embeddings and substantially greater alignment of image and text embeddings. Thus, the intra-modality separation leads to a better alignment and a substantial reduction of the modality gap. Furthermore, Figure 5(c) shows that AlignCLIP reduces the sparsity of the embeddings on the hypersphere.
(a) Original CLIP
(b) SharedCLIP
(c) AlignCLIP
Figure 5: DOSNES visualization of the multi-modal embeddings using CC3M
Comparison of Qualitative Examples. In Table 8, examples from the MSCOCO validation dataset is provided where the images convey the same general semantics but one of the ground truth texts provides more detailed information. On the left side of Table 8, two images with their corresponding ground truth captions are provided. We use CLIP, SharedCLIP and AlignCLIP for encoding the images and texts, and provide the cosine similarities on the right side of the table. As can be seen, when querying the first image, the similarity of the first image and the second text using the CLIP embeddings is higher in comparison to the ground truth caption. Suggesting that the second text will get selected as the predicted caption when using CLIP. This flaw still appears when using SharedCLIP for encoding the images and texts. However, when using AlignCLIP, the cosine similarity of the first image and the first text is higher in comparison to the second text, meaning that when querying the first image, the ground truth caption, which is semantically more correct in comparison to the second text, successfully gets selected. This suggests that the semantic regularization of the Intra-Modality Separation in AlignCLIP, which is calculated using the semantics of the text samples, potentially contributes in improving the retrieval performance.
CLIP
SharedCLIP
AlignCLIP
$T_{1}$ $T_{2}$ $T_{1}$ $T_{2}$ $T_{1}$ $T_{2}$
$I_{1}$ 0.37 0.42 0.53 0.55 0.61 0.58
$I_{2}$ 0.39 0.52 0.61 0.64 0.63 0.71
Table 8: Qualitative example of semantically similar samples from MSCOCO.
@@ 5 Related Work
Modality Gap in CLIP. The modality gap in CLIP’s embedding space was first studied by Liang et al. (2022). Authors showed that the modality gap is caused by a combination of the model initialization and the contrastive loss optimization. Furthermore, they showed that CLIP’s embedding space is very sparse such that the effective embedding space is an extremely narrow cone. In a similar study, Tyshchuk et al. (2023) measured the alignment of image and text embeddings in CLIP with a focus on the isotropic properties. More recently, Schrodi et al. (2024) showed that one of the key factors contributing in both the modality gap and the bias in CLIP is the information-imbalance between the two modalities, i.e., images often have much more detailed information of the scene in comparison to their corresponding text descriptions. Furthermore, Jiang et al. (2023) showed that even under the perfect alignment, the prediction error in downstream tasks cannot be smaller than the information gap that exists between the modalities. They further propose intra-modality as well as inter-modality regularization using augmented samples in order to improve the latent embeddings structures. In contrast to the previous work, we study the effectiveness of parameter sharing between the modality encoders as well as semantically-regularized separation of the uni-modal embeddings on reducing the modality gap, under the same amount of information gap across modalities.
@@ 6 Conclusion
This work investigates the potential of reducing the modality gap in CLIP by sharing the learnable parameter space of the vision and language encoders as well as enforcing a contrastive intra-modality separation objective. We further examine the effects of modality gap reduction by the aforementioned refinements in the performance of downstream tasks. Through extensive experiments, we show that SharedCLIP and AlignCLIP improve various zero-shot as well as fine-tuning downstream applications when compared to CLIP, while substantially improving the cross-modal alignment, and therefore, reducing the gap. Our work shows that it is possible to mitigate the modality gap in CLIP via parameter sharing and intra-modality separation without losing downstream performance.
@@ A.1 Training Dataset and Setup
We adopted a transformer encoder consisting of 12 layers and 12 heads in SharedCLIP and AlignCLIP. The same input pre-processing and augmentations employed in CLIP have been used for SharedCLIP as well as AlignCLIP, including random cropping of images to the size $224\times 224$. The image patch size for encoding visual data is set to $16\times 16$. When encoding texts, the maximum sequence length is set to 77 tokens and the vocabulary size for the embedding layer is set to 49,408. The output embedding dimension for both the vision and language modalities is set to 768. We chose our setup to be quite similar to that of the original CLIP with the ViT-B-16 backend, in order to ensure a fair comparison of CLIP and our proposed model. SBERT all-mpnet-base-v2 model has been utilized in AlignCLIP for re-scaling text–text and image–image cosine similarities is the .
@@
Model ImageNet-1K CIFAR-100 CIFAR-10 Flowers-102 Stanford Cars
Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
AlignCLIP (TT) 31.1 58.6 31.5 60.9 64.8 95.5 18.7 39.5 10 36.4
AlignCLIP (II) 32.8 60.6 36.5 66.4 69.3 97.8 18.8 40.3 11.8 38.1
AlignCLIP (II-TT) 32.4 60 31.2 61.8 66.4 96.9 18.7 39.9 11.8 37.1
Table 9: Accuracy scores for zero-shot image classification.
Model ImageNet-1K CIFAR-100 CIFAR-10 Flowers-102 Stanford Cars
AlignCLIP (TT) 51.2 64.5 86.3 74.3 41.6
AlignCLIP (II) 51.5 67.4 87.2 76.8 45.6
AlignCLIP (II-TT) 50.3 63.4 85.8 72.2 42.2
@@
A.2 Effects of text embedding separation
In Section 3.2, we proposed IMSep to enforce intra-modality separation among the image embeddings. A potential question that arises is that what happens if we enforce the separation amongst the text embeddings, or even, what happens of we enforce the separation inside the image embeddings as well as the text embeddings. To answer these questions, we trained versions of AlignCLIP without image-image separation and with only text-text separation, namely AlignCLIP-TT. In this case, we also used the re-scaling mechanism described in Section 3.2. Furthermore, we trained a version including both image-image embeddings separation and text-text embeddings separation, i.e., AlignCLIP-II-TT. We followed the same experimental setup described in Section A.1 and compare the results for zero-shot image classification as well as linear probing in Table 9 and Table 10, respectively.